# Rate-In: Information-Driven Adaptive Dropout Rates for Improved Inference-Time Uncertainty Estimation

### v3

@@
Rate-In: Information-Driven Adaptive Dropout Rates for Improved
Inference-Time Uncertainty Estimation
Tal Zeevi1, Ravid Shwartz-Ziv2, Yann LeCun2,3,
Lawrence H. Staib1†, John A. Onofrey1†
1Yale University, 2New York University, 3Meta FAIR
[…]
@@
{NoHyper}footnotetext: These authors contributed equally to this work.
@@ 1 Introduction
Figure 1: Rate-In maintains anatomically meaningful uncertainty estimates at high dropout rates, while standard methods degrade into noise. Left: Prostate MRI segmentation showing ground truth (yellow), model predictions (blue), and errors (red). Right: Uncertainty maps (brighter regions indicate higher uncertainty) at increasing dropout rates (1%-16%). While constant dropout produces diffuse uncertainty that obscures anatomical details, Rate-In precisely highlights challenging regions, particularly along tissue boundaries, preserving clinical utility of the uncertainty estimates.
@@
Monte Carlo Dropout [2] is a popular and practical technique for approximating predictive uncertainty in neural networks. By performing stochastic forward passes with dropout activated during inference, Monte Carlo Dropout samples from an approximate posterior distribution over the network’s weights, enabling the estimation of both the mean prediction and the associated uncertainty. This method offers a balance between computational efficiency and uncertainty estimation quality, making it suitable for large-scale applications.
A common limitation of Monte Carlo Dropout is the use of fixed dropout rates during inference, typically the same rates used during training. These static dropout rates are applied uniformly across all layers and inputs, regardless of the varying characteristics of individual inputs and the differing sensitivities of network layers. This rigidity can lead to suboptimal uncertainty estimates [3]: fixed dropout rates may either inject excessive noise, degrading the model’s predictive performance, or fail to introduce sufficient variability to capture the true predictive uncertainty.
As illustrated in Figure 1, the way we estimate this uncertainty can dramatically affect our ability to identify truly challenging regions in tasks such as medical image segmentation. Traditional uncertainty estimation methods often produce overly diffuse uncertainty patterns that fail to distinguish between genuinely difficult decisions and simpler ones. This limitation becomes particularly critical in medical applications, where highlighting truly uncertain regions can guide clinical attention to areas requiring additional scrutiny.
Existing approaches have explored optimizing dropout rates during training [4, 5, 6], often relying on labeled data to learn optimal dropout parameters. However, these methods result in fixed dropout rates that remain static during inference, unable to adapt to new or diverse input distributions encountered after deployment. Moreover, in many practical scenarios, especially in medical applications, ground truth labels may not be available during inference, necessitating methods that can adapt without supervision.
@@
In this paper, we introduce Rate-In, an algorithm that dynamically adjusts dropout rates during inference by quantifying the functional information loss in feature maps induced by dropout. Rate-In operates without requiring ground truth labels, making it suitable for deployment in real-world applications where labeled data is unavailable or scarce. By treating each layer’s dropout-induced information loss as a controllable parameter, Rate-In adaptively tunes dropout rates to maintain perceptual quality across diverse tasks and architectures.
@@
  • We demonstrate the effectiveness of Rate-In through extensive empirical evaluations on synthetic data and real-world medical imaging tasks, showing that it improves uncertainty estimation and calibration compared to fixed or heuristic dropout rates, without compromising predictive performance.
@@ 2 Background
In this section, we provide the necessary background to understand our proposed method. We begin by discussing the importance of uncertainty estimation in neural networks and the existing approaches to achieve it. We then go over the Monte Carlo Dropout as a practical method for uncertainty estimation, highlighting its limitations due to fixed dropout rates during inference. Finally, we introduce the application of information theory in deep learning, particularly how it can inform the adaptive adjustment of dropout rates to improve uncertainty estimation.
@@ 2.1 Uncertainty Estimation in Neural Networks
Accurately estimating both types of uncertainty is essential for building reliable models. Various methods have been proposed for uncertainty estimation in neural networks:
@@
  • Monte Carlo Dropout: Applies dropout at inference time to approximate Bayesian inference [2], providing a practical and computationally efficient method for uncertainty estimation.
2.2 Monte Carlo Dropout for Uncertainty Estimation
Monte Carlo (MC) Dropout leverages the dropout regularization technique [12] for uncertainty estimation by performing multiple stochastic forward passes through the network with dropout activated during inference [2]. This approach approximates a Bayesian posterior over the model parameters, allowing for the estimation of predictive uncertainty.
Given an input $\mathbf{x}$ and a neural network with parameters $\mathbf{W}$, the predictive distribution is approximated by averaging the predictions over $T$ stochastic forward passes:
$p(\mathbf{y}|\mathbf{x})\approx\frac{1}{T}\sum_{t=1}^{T}p(\mathbf{y}|\mathbf{x},\mathbf{W}_{t})$ (1)
@@ 2.3 Limitations of Fixed Dropout Rates
While MC Dropout is a practical method for uncertainty estimation, it commonly uses fixed dropout rates across all layers and inputs during inference [2]. This approach has limitations:
@@
  • Lack of Adaptability: Fixed rates do not account for the varying complexity of different inputs or the differing sensitivities of layers within the network. This rigidity limits the model’s ability to adapt to new or diverse data distributions.
These limitations highlight the need for methods that can adapt dropout rates dynamically during inference to improve uncertainty estimation without compromising performance.
@@ 2.4 Information Theory in Deep Learning
  • Information Bottleneck Principle: Proposes that layers in a neural network aim to compress input data while preserving relevant information for the output [13]. This principle has been used to analyze learning dynamics and generalization.
@@
However, most of these applications focus on the training phase of neural networks, using information theory to guide learning and improve representations. The potential of information theory to inform adaptive mechanisms during inference, particularly for uncertainty estimation, remains underexplored.
@@ 2.5 Applying Information Theory to Dropout
Interpreting dropout through the lens of information theory allows us to view it as introducing controlled noise into the network, effectively treating each layer as a noisy communication channel [7]. This perspective provides valuable insights into how dropout affects information flow within the network:
@@ 3 Rate-In: Adaptive Dropout at Inference
In this section, we present Rate-In, an algorithm that dynamically adjusts dropout rates during inference by quantifying the information loss induced in each layer’s feature maps. By viewing dropout as controlled noise injection and leveraging information-theoretic principles, Rate-In adapts dropout rates per layer and per input instance, enhancing uncertainty estimation without compromising predictive performance.
@@ 3.1 Reinterpreting Dropout as Noise Injection
Traditionally, dropout serves as a regularization technique that randomly drops units or connections during training [12, 18]. In the context of uncertainty estimation during inference, we reinterpret dropout as a form of noise injection into the neural signal. This perspective allows us to analyze dropout’s impact on information flow within the network.
By considering each layer as a communication channel [7], applying dropout introduces noise into this channel, reducing the mutual information between the layer’s input and output. The extent of information loss depends on the dropout rate and the characteristics of the activations in that layer.
@@ 3.2 Quantifying Information Loss
To adjust dropout rates dynamically, we need to quantify the information loss induced by dropout in each layer. We employ the concept of mutual information (MI), which measures the shared information between two random variables [7].
Given a layer with input activations $\mathbf{h}{\text{in}}$ and output activations $\mathbf{h}{\text{out}}$, the mutual information $I(\mathbf{h}{\text{in}};\mathbf{h}{\text{out}})$ quantifies how much information about $\mathbf{h}{\text{in}}$ is preserved in $\mathbf{h}{\text{out}}$. Dropout reduces this mutual information due to the introduced randomness.
@@ 3.3 The Rate-In Algorithm
Rate-In adjusts the dropout rate for each layer during inference by maintaining the information loss below a predefined threshold. The algorithm operates sequentially during the forward pass, adjusting dropout rates per layer and per input instance (Figure  2 and Algorithm  1).
@@
8:         Compute activations $\mathbf{h}^{(l)}{\text{in}}$ and $\mathbf{h}^{(l)}{\text{out}}$
9:         Estimate mutual information $I(\mathbf{h}^{(l)}{\text{in}};\mathbf{h}^{(l)}{\text{out}})$
10:         Compute information loss $\Delta I_{l}=I^{(l)}{\text{full}}-I(\mathbf{h}^{(l)}{\text{in}};\mathbf{h}^{(l)}{\text{out}})$
@@ 3.3.1 Algorithm Details
    Adjust Dropout Rate: If $\Delta I_{l}>\epsilon$, decrease $p_{l}$ to reduce information loss; otherwise, increase $p_{l}$ to introduce more variability.
@@
After adjusting dropout rates, perform a final forward pass with the adjusted rates to obtain $\hat{\mathbf{y}}$.
@@ 3.3.2 Implementation Considerations
Adjusting dropout rates can be formulated as an optimization problem. Simple strategies like gradient descent or rule-based adjustments (e.g., fixed step sizes) can be used to reach the desired information loss threshold.
@@
  • Share computations across similar inputs or batches.
@@ 3.4 Discussion of the Rate-In Approach
  • Adaptive Dropout Rates: By adjusting dropout rates per layer and per input instance, Rate-In accommodates the varying sensitivities of layers and the complexities of different inputs.
@@
  • No Need for Ground Truth Labels: Rate-In operates during inference without requiring labeled data, making it applicable in deployment scenarios where labels are unavailable.
@@
  • Computational Complexity: The need to estimate mutual information and adjust dropout rates can increase inference time.
  • Estimation Accuracy: Accurate estimation of mutual information is non-trivial, especially in high-dimensional spaces.
@@
In the next section, we present experimental results demonstrating Rate-In’s effectiveness across various tasks and architectures.
@@ 4 Experiments
In this section, we evaluate the effectiveness of the proposed Rate-In algorithm for predictive uncertainty estimation through MC simulations. We conduct experiments on both synthetic data and real-world medical imaging tasks, utilizing state-of-the-art pre-trained neural network architectures. Our evaluation focuses on three main objectives:
@@ 4.1 Experimental Setup
We evaluate our method on three categories of tasks:
@@
To validate our method in a controlled setting, we generate a one-dimensional regression dataset with varying levels of noise. The target variable is defined as $y=\sin(x)+\epsilon$, where $x$ is uniformly sampled from $[-3,3]$ and $\epsilon\sim\mathcal{N}(0,\sigma^{2})$. We create datasets with five different noise levels by setting $\sigma\in{0.1,0.2,0.3,0.4,0.5}$. Each dataset contains 100 training and test instances.
@@
For real-world classification tasks, we utilize three datasets from the MedMNIST v2 repository [19]:
@@
Baselines Methods
We compare Rate-In with the following methods:
@@ 4.1.1 Implementation Details
To directly compare Rate-In with classical dropout approaches, we set the information loss threshold $\epsilon$ equal to the fixed dropout rate $p^{(0)}$. $\delta$ was set to 0.01 based on preliminary experiments. The maximum number of iterations per layer was set to $N_{\text{max}}=30$.
For information loss estimation, we use the adaptive binning estimator [17], which efficiently approximates mutual information in high-dimensional feature spaces. Additionally, we measured the information using structural similarity index measure (SSIM) [22] (See Appendix B for more information and results). Monte Carlo simulation involves 30 forward passes per input, with dropout rates dynamically adjusted according to the measured information loss. We employ the following neural network architectures:
@@
  • Segmentation Tasks: nnU-Net [24], a U-Net-based architecture with dropout layers after each encoder and decoder block.
@@
For predictive performance, we use Mean Squared Error (MSE) for regression, Accuracy (ACC) for classification, and Dice Similarity Coefficient (DSC) for segmentation. To evaluate uncertainty estimation, we employ Expected Calibration Error (ECE)[25], Area Under the Accuracy-Rejection Curve (AUARC)[26], and Boundary Uncertainty Coverage (BUC) [27]. We also provide qualitative analyses through uncertainty maps and reliability diagrams.
Our code and complete implementation details are available in the supplementary material. The appendix includes full training protocols, hyperparameter settings, architectural details, and ablation studies.

### v4

@@
Abstract
@@ 1 Introduction
Figure 1: Rate-In maintains anatomically meaningful uncertainty estimates at high dropout rates, while standard methods degrade into noise. Left: Prostate MRI segmentation showing ground truth (yellow), model predictions (blue), and errors (red). Right: Uncertainty maps (brighter regions indicate higher uncertainty) at increasing dropout rates (1%-16%). While constant dropout produces diffuse uncertainty that obscures anatomical details, Rate-In precisely highlights challenging regions, particularly along tissue boundaries, preserving the clinical utility of uncertainty estimates.
@@
Monte Carlo (MC) Dropout [12] is a practical technique for approximating predictive uncertainty in neural networks. By running stochastic forward passes with dropout activated during inference, MC Dropout samples from an approximate posterior distribution over the network’s weights, enabling estimation of the mean prediction and the associated uncertainty. This method offers a balance between computational efficiency and uncertainty estimation quality, making it suitable for large-scale applications.
A common limitation of MC Dropout is the use of fixed dropout rates during inference, typically the same rates used during training. These static dropout rates are applied uniformly across all layers and inputs, regardless of the varying characteristics of individual inputs and the differing sensitivities of network layers. This rigidity can lead to suboptimal uncertainty estimates [33]: fixed dropout rates may either inject excessive noise, degrading the model’s predictive performance, or fail to introduce sufficient variability to capture the true predictive uncertainty.
As shown in Figure 1, how uncertainty is estimated can dramatically affect our ability to identify challenging regions in tasks such as medical image segmentation. Traditional uncertainty estimation methods tend to produce overly diffuse uncertainty patterns, failing to separate genuinely difficult decisions from simpler ones. This limitation becomes particularly critical in medical applications, where highlighting truly uncertain regions can guide clinical attention to areas requiring additional scrutiny.
Existing approaches have explored optimizing dropout rates during training [3, 13, 22], often relying on labeled data to learn optimal dropout parameters. However, these methods result in fixed dropout rates that remain static during inference, unable to adapt to new or diverse input distributions encountered after deployment. When the model’s architecture and parameters are inaccessible at inference, or the training dropout setting is unknown, post-hoc dropout methods enable uncertainty estimation without modifying training. Techniques like dropout injection at test time [26] and infer-dropout [29] apply dropout retrospectively to pre-trained models, offering flexibility in inference, such as widening predictive intervals in safety-critical settings. However, their reliance on fixed dropout rates makes them insensitive to data variability and may be suboptimal in heterogeneous settings like medical applications.
@@
In this paper, we introduce Rate-In, an algorithm that dynamically adjusts dropout rates during inference by quantifying the functional information loss in feature maps induced by dropout. Rate-In requires no ground truth labels, making it suitable for deployment in real-world applications where labeled data is unavailable or scarce. By treating each layer’s dropout-induced information loss as a controllable parameter, Rate-In adaptively tunes dropout rates to preserve perceptual quality across tasks and architectures.
@@
  • We demonstrate Rate-In’s effectiveness through extensive empirical evaluations on synthetic data and real-world medical imaging tasks, showing improved uncertainty estimation and calibration over fixed or heuristic dropout rates, without compromising predictive performance.
@@ 2 Background
This section provides the necessary background for our method. We first discuss the role of uncertainty estimation in neural networks and the existing approaches to achieve it. Next, we examine Monte Carlo (MC) Dropout as a practical method for uncertainty estimation, highlighting its limitations due to fixed dropout rates during inference. Finally, we introduce the application of information theory in deep learning, particularly how it can inform the adaptive adjustment of dropout rates to improve uncertainty estimation.
@@ 2.1 Uncertainty Estimation in Neural Networks
Accurately estimating both types of uncertainty is essential for building reliable models. Various methods have been proposed for uncertainty estimation in neural networks; below are key representative approaches:
@@
  • Test-time Data Augmentation: Repeated input perturbation during inference (e.g., rotation, blurring) to approximate a predictive distribution. Effective in domains where transformations are guided by domain knowledge [29].
  • Noise Injection Techniques: Introduces controlled noise to the model (e.g., Gaussian noise added to weights) to assess sensitivity beyond input-space augmentation [29].
  • Monte-Carlo (MC) Techniques: Use stochastic sampling to approximate Bayesian uncertainty. MC Dropout applies dropout at inference to sample the model’s parameter space [12]. MC Batch Normalization perturbs batch statistics by sampling batches from the training data during inference, making it effective when training data is accessible [39]. Masksembles employs fixed binary masks with controlled sparsity and overlap, balancing MC Dropout’s randomness with Deep Ensembles’ independence, but without posterior sampling [10]. MC Layer Normalization introduces stochasticity by subsampling features (e.g., via fixed-rate dropout) and computing layer norm statistics on the sampled feature space [11].
2.2 MC Dropout for Uncertainty Estimation
Monte Carlo Dropout leverages the dropout regularization technique [38] for uncertainty estimation by performing multiple stochastic forward passes through the network with dropout activated during inference [12]. This approach approximates a Bayesian posterior over the model parameters, allowing for the estimation of predictive uncertainty.
Given an input $\mathbf{x}$ and a neural network with parameters $\mathbf{W}$, the predictive distribution $\mathbf{y}$ is approximated by averaging the predictions over $T$ stochastic forward passes:
$p(\mathbf{y}|\mathbf{x})\approx\frac{1}{T}\sum_{t=1}^{T}p(\mathbf{y}|\mathbf{x},\mathbf{W}_{t}),$ (1)
@@ 2.3 Limitations of Fixed Dropout Rates
While MC Dropout is a practical method for uncertainty estimation, it typically applies fixed dropout rates across all layers and inputs during inference [12]. This approach has limitations:
@@
  • Lack of Adaptability: Fixed rates ignore the varying complexity of inputs or the differing sensitivities of layers within the network. This rigidity limits the model’s ability to adapt to new or diverse data distributions.
These limitations highlight the need for methods that adapt dropout rates dynamically during inference to improve uncertainty estimation without compromising performance.
@@ 2.4 Information Theory in Deep Learning
  • Information Bottleneck Principle: Proposes that neural network layers aim to compress input data while preserving relevant information for the output [37]. It has been used to analyze learning dynamics and generalization.
@@
However, most of these applications focus on the training of neural networks, using information theory to guide learning and improve representations. The potential of information theory to drive adaptive mechanisms during inference, particularly for uncertainty estimation, remains underexplored.
@@ 2.5 Applying Information Theory to Dropout
Interpreting dropout through information theory allows us to view it as introducing controlled noise into the network, effectively treating each layer as a noisy communication channel [8]. This perspective provides valuable insights into dropout’s impact information flow within the network:
@@ 3 Rate-In: Adaptive Dropout at Inference
In this section, we present Rate-In, an algorithm that dynamically adjusts dropout rates during inference by quantifying the information loss in each layer’s feature maps. Treating dropout as controlled noise injection and leveraging information-theoretic principles, Rate-In adapts dropout rates per layer and per input instance, enhancing uncertainty estimation without compromising predictive performance.
@@ 3.1 Reinterpreting Dropout as Noise Injection
Traditionally, dropout serves as a regularization technique that randomly drops units or connections during training [38, 40]. For uncertainty estimation during inference, we reinterpret dropout as a form of noise injection into the neural signal. This perspective allows us to analyze dropout’s impact on information flow within the network.
Treating each layer as a communication channel [8], dropout introduces noise into this channel, reducing the mutual information between the layer’s input and output. The extent of information loss depends on the dropout rate and the characteristics of the activations in that layer.
@@ 3.2 Quantifying Information Loss
To dynamically adjust dropout rates, we need to quantify the information loss induced by dropout in each layer. We use mutual information (MI), which measures the shared information between two random variables [8].
Given a layer with input activations $\mathbf{h}_{\text{in}}$ and output activations $\mathbf{h}_{\text{out}}$, the mutual information $I(\mathbf{h}_{\text{in}};\mathbf{h}_{\text{out}})$ quantifies how much information about $\mathbf{h}_{\text{in}}$ is preserved in $\mathbf{h}_{\text{out}}$. Dropout reduces this mutual information due to the introduced randomness.
@@ 3.3 The Rate-In Algorithm
Rate-In adjusts the dropout rate for each layer during inference by maintaining the information loss below a predefined threshold. The algorithm operates sequentially during the forward pass, adjusting dropout rates per layer and per input instance (Figure 2 and Algorithm 1).
@@ 3.3.1 Algorithm Details
8:         Compute activations $\mathbf{h}^{(l)}_{\text{in}}$ and $\mathbf{h}^{(l)}_{\text{out}}$
9:         Estimate mutual information $I(\mathbf{h}^{(l)}_{\text{in}};\mathbf{h}^{(l)}_{\text{out}})$
10:         Compute information loss $\Delta I_{l}=I^{(l)}_{\text{full}}-I(\mathbf{h}^{(l)}_{\text{in}};\mathbf{h}^{(l)}_{\text{out}})$
@@
    Adjust Dropout Rate: If $\Delta I_{l}>\epsilon$, decrease $p_{l}$ to reduce information loss; else, increase $p_{l}$ to enhance variability.
@@
After adjusting dropout rates for all layers, perform a final forward pass with adjusted rates to obtain $\hat{\mathbf{y}}$.
@@ 3.3.2 Implementation Considerations
Adjusting dropout rates can be formulated as an optimization problem. Simple strategies like gradient descent or rule-based updates (e.g., fixed step sizes) can achieve the target information loss threshold.
@@
  • Share computations across similar inputs or batchs.
@@ 3.4 Discussion of the Rate-In Approach
  • Adaptive Dropout Rates: By adjusting dropout rates per layer and input instance, Rate-In accommodates varying layers sensitivities and input complexities.
@@
  • No Need for Ground Truth Labels: Rate-In operates during inference without labeled data, making it applicable in deployment scenarios where labels are unavailable.
@@
  • Computational Complexity: The need to estimate MI and adjust dropout rates can increase inference time.
  • Estimation Accuracy: Accurately estimating MI is non-trivial, especially in high-dimensional spaces.
@@
Next, we present experimental results demonstrating Rate-In’s effectiveness across various tasks and architectures.
@@ 4 Experiments
In this section, we evaluate the effectiveness of the proposed Rate-In algorithm for predictive uncertainty estimation through Monte Carlo (MC) simulations. We test it on synthetic data and real-world medical imaging tasks with state-of-the-art pre-trained neural network architectures. Our evaluation focuses on three main objectives:
@@ 4.1 Experimental Setup
We evaluate our method on four categories of tasks:
@@
To validate our method in a controlled setting, we generate a one-dimensional regression dataset with varying levels of noise. The target variable is defined as $y=\sin(x)+\epsilon$, where $x$ is uniformly sampled from $[-3,3]$ and $\epsilon\sim\mathcal{N}(0,\sigma^{2})$. We create datasets with five different noise levels by setting $\sigma\in\{0.1,0.2,0.3,0.4,0.5\}$. Each dataset contains 100 training and test instances.
@@
For real-world classification tasks, we use datasets from MedMNIST v2 [42]:
@@
Out of Distribution Robustness
To evaluate robustness to real-world medical imaging artifacts, we used MedMNIST-C [9], a dataset with controlled task- and modality-specific corruptions, similar to ImageNet-C [18]. These include common artifacts such as air bubbles, defocus, and motion blur, applied at varying severity levels. We tested the classification datasets in their corrupted versions: PathMNIST-C, BloodMNIST-C, and TissueMNIST-C.
We compare Rate-In with the following baseline methods:
@@
Formal definitions of these dropout strategies are provided in Appendix A Section  6.2.3.
@@ 4.1.1 Implementation Details
To ensure a fair comparison, Rate-In’s information loss threshold ($\epsilon$) was set equal to the dropout rate ($p^{(0)}$) of other benchmarks, emphasizing the conversion of impulse noise to contextual noise. With $p_{0}=p$, all methods start equally, but Rate-In dynamically adjusts dropout rates based on layer-wise redundancy. $\delta$ was set to 0.01 and $N_{\text{max}}$ to 30 based on preliminary experiments.
For information loss estimation, we use the adaptive binning estimator [7] to efficiently approximate mutual information in high-dimensional feature spaces. We also assess feature preservation using the structural similarity index (SSIM) [41] (See Appendix B for details and results). MC simulations use 30 forward passes per input, with dropout rates dynamically adjusted based on measured information loss. We employ the following network architectures:
@@
  • Segmentation Tasks: nnU-Net [20], a U-Net-based architecture with dropout after each block.
@@
For predictive performance, we use Mean Squared Error (MSE) for regression, Accuracy (ACC) for classification, and Dice Similarity Coefficient (DSC) for segmentation. For uncertainty estimation, we employ Expected Calibration Error (ECE) [16], Area Under the Accuracy-Rejection Curve (AUARC) [15], and Boundary Uncertainty Coverage (BUC) [43]. We also provide qualitative analyses through uncertainty maps. Full metric formulations are in Appendix A Section  6.2.4.
Computational Complexity
We analyze the theoretical complexity of Rate-In optimization and validate it empirically. Sensitivity to key parameters - initial dropout rate, noise level, information loss threshold, and dataset size, is studied on synthetic data, while inference overhead is assessed in real-world classification and segmentation tasks.
Reproducibility and Ablation Studies
Code and implementation details are in the supplementary material. Appendix A covers methodology, training protocols, and evaluation. Appendix B includes additional results, ablations, and complexity analysis. Appendix C discusses Rate-In’s theoretical properties and implications.